OpenStudy (anonymous):

In ∆ABC, if sin A = 4/5 and tan A = 4/3 , then what is cos A?

OpenStudy (gorv):

sin= perpendicular/hypotenuse=4/5

OpenStudy (gorv):

perpendicular=4 hypotenuse=5

OpenStudy (anonymous):

Thank you! :)

OpenStudy (gorv):

tan=perpendicular /base=4/3 bas3=3

OpenStudy (gorv):

cos= base /hypotenuse

OpenStudy (gorv):

put values